A bag contains 3 biased coins B1, B2 AND B3 whose probability of falling head wise are 1/3, 2/3 and 3/4 respectively. A coin is drawn at random and tossed, fell head wise. Find the probability that same coin is tossed again will fall head wise.

Grade:12th pass

2 Answers

Aman
3 years ago
P(b1)=p(b2)=p(b3)=1/3
E: randomly selected coin falls headwise
P(e/b1)=4/21
Pe/b2)=8/21
P(e/b3)=9/21
4/21*1/3+8/21*1/3+9/21*3/4=23/36
 
 
Answer=23/36
Vikas TU
3 years ago
Dear student 
This is a Bayes rule problem.
P(Biased|Heads)=P(Heads|Biased)P(Biased)/{P(Heads|Biased)P(Biased)+P(Heads|Fair)P(Fair)}.
P(Heads|Biased)=4/5 and I'll leave the rest to you.